Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
Answer by Alan3354(69443)

N=number of nickels
D=number of dimes
So 5N+10D=565
-------------
If:
10(D*2)+5(N+8) = 1045, or 20D+5N+40 = 1045
2 equations in 2 unknowns.
5N+10D=565 eqn 1
5N+20D=1005 eqn 2
------------------
Subtract eqn 2 from eqn 1
-10D = -440
D = 44, the number of dimes
Sub D=44 into either eqn, and find that N=25
-----------------
Joe must be VERY proud to have 44 dimes and 25 nickels. Jeez.